# KaSA: Knowledge-Aware Singular-Value Adaptation of Large Language Models

### v1

@@ 1 Introduction
  • KaSA features a linear framework that allows seamless integration of the singular value adaptation module with the SVD truncated model architecture, inducing no inference latency. Our method also supports training distinct adaptation modules for different tasks, all sharing a single base model, thereby reducing the storage needs for task-switching.
  • We conduct extensive experiments on NLU, NLG, instruction following, and commonsense reasoning tasks using popular LLMs on well-known benchmarks. Our method consistently outperforms FFT and 14 PEFT baselines across different benchmarks and synthetic datasets, demonstrating its efficacy and adaptability.
@@ 4 Experiments
In this section, we evaluate KaSA’s efficacy across different downstream tasks, including natural language understanding (NLU), natural language generation (NLG), instruction following, and commonsense reasoning. For NLU tasks, we evaluate KaSA with RoBERTa (Liu et al., 2021) and DeBERTaV3 (He et al., 2022b) on the GLUE (Wang et al., 2018) benchmark. For NLG tasks, we assess our method with GPT-2 (Radford et al., 2019) on the E2E NLG Challenge (Novikova et al., 2017) benchmark. We further assess instruction following performance with popular LLMs, including LLaMA3 8B (Meta, 2024), Mistal 7B (Jiang et al., 2023), Gemma 7B (Gemma Team, 2024), and LLaMA2 13B (Touvron et al., 2023b). These models are fine-tuned with different PEFT methods using four synthetic datasets generated by GPT4o, each tailored to summarization, classification, coding, and closed QA. GPT4o is then employed as a judge to evaluate the fine-tuned models’ performance, assigning scores on a scale of 10. We also follow (Kopiczko et al., 2023) and (Gao et al., 2024) to fine-tune the four models on the Alpaca dataset (Taori et al., 2023b) and report evaluation results on MT-Bench, with GPT4 serving as the judge, yielding scores within 10. Additionally, we substantiate KaSA’s generality by fine-tuning LLaMA2 7B and LLaMA3 8B models on the Commonsense170K dataset (Hu et al., 2023), which includes training sets from eight commonsense reasoning datasets, and evaluating them on individual test sets of these constituent datasets. Finally, we conduct ablation studies to investigate the impacts of different components, budget parameter scalability, and the distribution of knowledge-aware singular values across various layers. All experiments are conducted on NVIDIA A100-SXM4 (80GB) GPUs, except for the NLU experiments, which are conducted on NVIDIA GeForce RTX 3090 (24GB) GPUs.
@@ 4.1 Baselines
$\bullet$ SVD-based methods Considering that our method is associated with SVD, we chose other SVD-based PEFT baselines: 1) AdaLoRA (Zhang et al., 2022); 2) PiSSA (Meng et al., 2024); 3) MiLoRA (Wang et al., 2024a); 4) SARA (Gu et al., 2024); and 5) CorDA (Yang et al., 2024).
@@ 4.2 Natural Language Understanding
Main Results. Table 1 presents the performance of RoBERTa-base and RoBERTa-large models fine-tuned using our KaSA in contrast to PEFT baselines. KaSA achieves the best performance across all datasets except MRPC and STS-B for the RoBERTa-base model. Notably, KaSA registers the highest average performances for both RoBERTa models: 86.3% for RoBERTa-base and 89.0% for RoBERTa-large. This underscores the effectiveness, adaptability, and scalability of our proposed approach. In a significant comparison with FFT, our KaSA, which utilizes merely up to 0.24% (approximately 0.3M/125.0M) of trainable parameters, outperforms FFT in 13 out of 14 scenarios and matches its performance on the STS-B dataset for the RoBERTa-base model. Furthermore, as demonstrated in Table 2, the DeBERTaV3-base results consistently surpass all baseline performances across the datasets, with the exception of STS-B, achieving the highest average performance of 88.72%. This further validates the efficacy of our method across different model architectures.
4.3 Natural Language Generation
Models and Datasets. For NLG tasks, we employ KaSA and other PEFT baselines to fine-tune both GPT-2 Medium (355M) and GPT-2 Large (774M) models (Radford et al., 2019) on the well-established E2E (End-to-End) NLG Challenge benchmark (Novikova et al., 2017), which focuses on restaurant domain information. The statistics of the E2E NLG Challenge benchmark and the evaluation metrics applied are detailed in Appendix C.2.
Implementation Details. We adopt the experimental configurations delineated in (Hu et al., 2021; Gu et al., 2024) for the fine-tuning of query and value weights within each Transformer block, setting a rank of $r=4$. The AdamW optimizer is employed, paired with a linear learning rate schedule over 5 epochs. The reported results represent the mean outcomes from 3 runs, each initialized with a distinct random seed, selecting the performance at the last epoch of each run for comparison. For further details on the hyperparameters utilized, refer to Appendix E.2.
Main Results.
Model(Method) # Trainable Parameters BLEU NIST METEOR ROUGE-L CIDEr
[…]
@@ 4.4 Instruction Following
Table 4: Instruction following evaluation results with average scores for the most popular LLMs fine-tuned on the 128k synthetic datasets and the Alpaca dataset, and evaluated by GPT4o and GPT4 with the scores within 10 on test subsets and MT-Bench, respectively. Significance is tested at the $\alpha=0.05$ level, and the p-value of the significance tests is reported.
@@
LoRA 3.21M 5.98$\pm$0.3 (p=0.001) 7.29$\pm$0.2 (p=0.002) 7.75$\pm$0.2 (p=0.049) 8.18$\pm$0.2 (p=0.002) 4.32$\pm$0.4 (p=0.032)
PiSSA 3.21M 6.23$\pm$0.2 (p=0.002) 7.88$\pm$0.1 (p=0.730) 7.80$\pm$0.1 (p=0.018) 8.22$\pm$0.2 (p=0.003) 4.66$\pm$0.3 (p=0.182)
MiLoRA 3.21M 6.30$\pm$0.1 (p=0.001) 7.62$\pm$0.2 (p=0.067) 7.71$\pm$0.2 (p=0.028) 8.27$\pm$0.3 (p=0.029) 4.53$\pm$0.2 (p=0.041)
KaSA 3.22M 6.88$\pm$0.2 7.92$\pm$0.2 8.01$\pm$0.1 8.69$\pm$0.1 4.97$\pm$0.3
@@
LoRA 3.40M 5.07$\pm$0.3 (p=0.007) 5.72$\pm$0.2 (p=0.000) 6.17$\pm$0.4 (p=0.034) 7.39$\pm$0.2 (p=0.034) 4.18$\pm$0.3 (p=0.057)
PiSSA 3.40M 5.46$\pm$0.2 (p=0.103) 5.86$\pm$0.3 (p=0.002) 6.41$\pm$0.2 (p=0.048) 7.24$\pm$0.2 (p=0.007) 4.24$\pm$0.2 (p=0.043)
MiLoRA 3.40M 5.33$\pm$0.2 (p=0.025) 5.89$\pm$0.4 (p=0.006) 6.52$\pm$0.2 (p=0.158) 7.28$\pm$0.3 (p=0.031) 4.29$\pm$0.2 (p=0.074)
KaSA 3.41M 5.72$\pm$0.2 6.82$\pm$0.3 6.74$\pm$0.2 7.75$\pm$0.2 4.58$\pm$0.2
@@
LoRA 3.40M 6.12$\pm$0.3 (p=0.019) 7.20$\pm$0.4 (p=0.016) 7.37$\pm$0.2 (p=0.006) 6.02$\pm$0.2 (p=0.002) 4.19$\pm$0.3 (p=0.020)
PiSSA 3.40M 6.35$\pm$0.1 (p=0.028) 7.31$\pm$0.3 (p=0.011) 7.59$\pm$0.1 (p=0.028) 6.18$\pm$0.3 (p=0.018) 4.26$\pm$0.2 (p=0.013)
MiLoRA 3.40M 6.37$\pm$0.2 (p=0.083) 7.61$\pm$0.1 (p=0.014) 7.65$\pm$0.2 (p=0.128) 6.39$\pm$0.1 (p=0.029) 4.32$\pm$0.2 (p=0.025)
KaSA 3.41M 6.55$\pm$0.2 7.83$\pm$0.1 7.89$\pm$0.2 6.81$\pm$0.3 4.71$\pm$0.2
@@
LoRA 6.55M 6.23$\pm$0.4 (p=0.023) 7.38$\pm$0.2 (p=0.005) 7.54$\pm$0.2 (p=0.005) 6.25$\pm$0.3 (p=0.001) 4.43$\pm$0.3 (p=0.011)
PiSSA 6.55M 6.47$\pm$0.3 (p=0.062) 7.45$\pm$0.3 (p=0.031) 7.83$\pm$0.1 (p=0.049) 6.54$\pm$0.3 (p=0.006) 4.39$\pm$0.2 (p=0.001)
MiLoRA 6.55M 6.45$\pm$0.2 (p=0.020) 7.63$\pm$0.1 (p=0.032) 7.85$\pm$0.1 (p=0.064) 6.82$\pm$0.2 (p=0.028) 4.51$\pm$0.3 (p=0.024)
KaSA 6.56M 6.86$\pm$0.2 7.92$\pm$0.2 8.09$\pm$0.2 7.12$\pm$0.1 4.95$\pm$0.1
@@
Main Results. In Table 4, the results show that KaSA consistently surpasses LoRA, PiSSA, and MiLoRA across four 128k synthetic datasets, regardless of the model used. Notably, Gemma 7B and LLaMA3 8B, fine-tuned with KaSA, even surpass FFT in the classification, summarization, and coding datasets. In the evaluation using MT-Bench, KaSA consistently outperforms FFT and PEFT baselines on all models, showing remarkable efficacy. With p-values less than 0.05 in 9 out of 12 experimental settings on MT-Bench, KaSA demonstrates significant performance improvements over LoRA, PiSSA, and MiLoRA. These results further highlight the effectiveness, robustness, and adaptability of our method.
@@ 5 Conclusion
In this paper, we introduce a PEFT method, KaSA, which incorporates SVD with knowledge-aware singular values for dynamic knowledge activation of parametric knowledge according to their relevance to a given task. KaSA commences knowledge-based SVD truncation of minor singular value components to remove noisy knowledge within the base model. Subsequently, it reparameterizes task-specific updates in the SVD form, leveraging knowledge-aware singular values for dynamic knowledge activation according to relevance. Our extensive experiments on various LLMs across tasks in NLU, NLG, instruction following, and commonsense reasoning reveal that KaSA consistently surpasses FFT and a variety of popular PEFT baselines across well-known benchmarks and our synthetic datasets, highlighting the superiority of our method.
@@ Appendix A Pseudocode for KaSA
21 self.delta_sigma = torch.diag(nn.Parameter(torch.randn(self.r), requires_grad=True))
@@
26 Delta_W = self.delta_u @ self.delta_sigma @ self.delta_v
@@ Appendix F Additional Experimental Results
F.1 Components Ablation Study on SST-2, QNLI, and STS-B
@@
F.2 Rank $k$ of Knowledge-Based SVD Truncation
@@
F.3 Rank $r$ of Knowledge-Aware Singular-Value Adaptation
@@
F.4 Parameter Initialization of $\Delta\mathbf{W}=\Delta\mathbf{U}\Delta\mathbf{\Sigma}\Delta\mathbf{V^{\top}}$
@@
F.5 Singular-Value and Orthogonal Regularization
@@
F.6 Hyperparameter Sensitivity Analysis
@@
F.7 Efficiency and Complexity Analysis
@@ G.1 Initialization Dilemmas of $\Delta\mathbf{W}$ in PiSSA and MiLoRA
$\Delta\mathbf{W}=\mathbf{A}\mathbf{B}=\mathbf{U}_{pri}\mathbf{\Sigma}_{pri}{\mathbf{V}}_{pri}^{\top},\ \mathbf{W}_{pri}\rightarrow\Delta\mathbf{W}$ (19)
@@
$\Delta\mathbf{W}$ $=\mathbf{A^{\prime}}\mathbf{B^{\prime}}=\mathbf{U^{\prime}}_{res}\mathbf{\Sigma^{\prime}}_{res}{\mathbf{V^{\prime}}}_{res}^{\top},\ \mathbf{W^{\prime}}_{res}\rightarrow\Delta\mathbf{W}$ (23)

### v2

@@ 1 Introduction
  • Our KaSA features a linear framework that allows seamless integration of the singular value adaptation module with the SVD truncated model architecture, inducing no inference latency. Furthermore, our method supports training distinct adaptation modules for different tasks, all sharing a single base model, thereby reducing the storage needs for task-switching.
  • We conduct extensive experiments on NLU, NLG, instruction following, and commonsense reasoning tasks using popular LLMs on well-known benchmarks. Our KaSA consistently outperforms FFT and 14 PEFT baselines across different benchmarks and synthetic datasets, demonstrating its efficacy and adaptability.
@@ 4 Experiments
In this section, we evaluate KaSA’s efficacy across different downstream tasks, including natural language understanding (NLU), natural language generation (NLG) (see Appendix F.2), instruction following, and commonsense reasoning. For NLU tasks, we evaluate KaSA with RoBERTa (Liu et al., 2021) and DeBERTaV3 (He et al., 2022b) on the GLUE (Wang et al., 2018) benchmark. For NLG tasks, we assess our method with GPT-2 (Radford et al., 2019) on the E2E NLG Challenge (Novikova et al., 2017) benchmark. We further assess instruction following performance with popular LLMs, including LLaMA3 8B (Meta, 2024), Mistal 7B (Jiang et al., 2023), Gemma 7B (Gemma Team, 2024), and LLaMA2 13B (Touvron et al., 2023b). These models are fine-tuned with different PEFT methods using four synthetic datasets generated by GPT4o, each tailored to summarization, classification, coding, and closed QA. GPT4o is then employed as a judge to evaluate the fine-tuned models’ performance, assigning scores on a scale of 10. We also follow (Kopiczko et al., 2023) and (Gao et al., 2024) to fine-tune the four models on the Alpaca dataset (Taori et al., 2023b) and report evaluation results on MT-Bench, with GPT4 serving as the judge, yielding scores within 10. Additionally, we substantiate KaSA’s generality by fine-tuning LLaMA2 7B and LLaMA3 8B models on the Commonsense170K dataset (Hu et al., 2023), which includes training sets from eight commonsense reasoning datasets, and evaluating them on individual test sets of these constituent datasets. Finally, we conduct ablation studies to investigate the impacts of different components, budget parameter scalability, and the distribution of knowledge-aware singular values across various layers. All experiments are conducted on NVIDIA A100-SXM4 (80GB) GPUs, except for the NLU experiments, which are conducted on NVIDIA GeForce RTX 3090 (24GB) GPUs.
@@ 4.1 Baselines
$\bullet$ SVD-based methods Considering that our method is associated with SVD, we chose SVD-based PEFT baselines: 1) AdaLoRA (Zhang et al., 2022); 2) PiSSA (Meng et al., 2024); 3) MiLoRA (Wang et al., 2024a); 4) SARA (Gu et al., 2024); and 5) CorDA (Yang et al., 2024).
@@ 4.2 Natural Language Understanding
Main Results. Table 1 presents the performance of RoBERTa-base and RoBERTa-large models fine-tuned using our KaSA in contrast to PEFT baselines. KaSA achieves the best performance across all datasets except MRPC and STS-B for the RoBERTa-base model. Notably, KaSA registers the highest average performances for both RoBERTa models: 86.3% for RoBERTa-base and 89.0% for RoBERTa-large. This underscores the effectiveness, adaptability, and scalability of our proposed approach. In a significant comparison with FFT, our KaSA, which utilizes merely up to 0.24% (approximately 0.3M/125.0M) of trainable parameters, outperforms FFT in 13 out of 14 scenarios and matches its performance on the STS-B dataset for the RoBERTa-base model. The results from DeBERTaV3-base are presented in Appendix F.1.
@@ 4.3 Instruction Following
Table 2: Instruction following evaluation results with average scores for the most popular LLMs fine-tuned on the 128k synthetic datasets and the Alpaca dataset, and evaluated by GPT4o and GPT4 with the scores within 10 on test subsets and MT-Bench, respectively.
@@
LoRA 3.21M 5.98 7.29 7.75 8.18 4.32
PiSSA 3.21M 6.23 7.88 7.80 8.22 4.66
MiLoRA 3.21M 6.30 7.62 7.71 8.27 4.53
KaSA 3.22M 6.88 7.92 8.01 8.69 4.97
@@
LoRA 3.40M 5.07 5.72 6.17 7.39 4.18
PiSSA 3.40M 5.46 5.86 6.41 7.24 4.24
MiLoRA 3.40M 5.33 5.89 6.52 7.28 4.29
KaSA 3.41M 5.72 6.82 6.74 7.75 4.58
@@
LoRA 3.40M 6.12 7.20 7.37 6.02 4.19
PiSSA 3.40M 6.35 7.31 7.59 6.18 4.26
MiLoRA 3.40M 6.37 7.61 7.65 6.39 4.32
KaSA 3.41M 6.55 7.83 7.89 6.81 4.71
@@
LoRA 6.55M 6.23 7.38 7.54 6.25 4.43
PiSSA 6.55M 6.47 7.45 7.83 6.54 4.39
MiLoRA 6.55M 6.45 7.63 7.85 6.82 4.51
KaSA 6.56M 6.86 7.92 8.09 7.12 4.95
@@
Main Results. In Table 2, the results show that KaSA consistently surpasses LoRA, PiSSA, and MiLoRA across four 128k synthetic datasets, regardless of the model used. Notably, Gemma 7B and LLaMA3 8B, fine-tuned with KaSA, even surpass FFT in the classification, summarization, and coding datasets. In the evaluation using MT-Bench, KaSA consistently outperforms FFT and PEFT baselines on all models, showing remarkable efficacy. With significance tests showing ($p<0.05$) in 9 out of 12 experimental settings on MT-Bench, KaSA demonstrates significant performance improvements over LoRA, PiSSA, and MiLoRA. These results further highlight the effectiveness, robustness, and adaptability of our method.
@@ 5 Conclusion
In this paper, we introduce a PEFT method, KaSA, which incorporates SVD with knowledge-aware singular values for dynamic activation of parametric knowledge according to their relevance to the given tasks. KaSA commences with knowledge-based SVD truncation of minor singular value components to remove noisy knowledge within the base model. Subsequently, it reparameterizes task-specific updates in the SVD form, leveraging knowledge-aware singular values for dynamic knowledge activation according to relevance. Our extensive experiments on various LLMs across tasks in NLU, NLG, instruction following, and commonsense reasoning reveal that KaSA consistently surpasses FFT and a variety of prevailing PEFT baselines across well-known benchmarks and our synthetic datasets, highlighting the superiority of our method.
Acknowledgments
Jing Tang’s work was partially supported by National Key R&D Program of China under Grant No. 2023YFF0725100 and No. 2024YFA1012701, by the National Natural Science Foundation of China (NSFC) under Grant No. 62402410 and No. U22B2060, by Guangdong Provincial Project (No. 2023QN10X025), by Guangdong Basic and Applied Basic Research Foundation under Grant No. 2023A1515110131, by Guangzhou Municipal Science and Technology Bureau under Grant No. 2023A03J0667 and No. 2024A04J4454, by Guangzhou Municipal Education Bureau (No. 2024312263), and by Guangzhou Municipality Big Data Intelligence Key Lab (No. 2023A03J0012), Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628) and Guangzhou Municipal Key Laboratory of Financial Technology Cutting-Edge Research (No. 2024A03J0630). This work was also supported by IITP grant funded by the Korea government(MSIT)[RS-2023-00215959, Development of Access Agnostic wired and wireless integrated optical access technology].
@@ Appendix A Pseudocode for KaSA
21 self.delta_sigma = nn.Parameter(torch.randn(self.r), requires_grad=True)
@@
26 Delta_W = self.delta_u @ torch.diag(self.delta_sigma) @ self.delta_v
@@ Appendix F Additional Experimental Results
F.1 Natural Language Understanding on DeBERTaV3-base
(Acc.) (Acc.) (Mcc.) (Acc.) (Acc.) (Pcc.) Avg.
As demonstrated in Table 11, the DeBERTaV3-base results consistently surpass all baseline performances across the datasets, with the exception of STS-B, achieving the highest average performance of 88.72%. This further validates the efficacy of our method across different model architectures.
F.2 Natural Language Generation
Models and Datasets. For NLG tasks, we employ KaSA and other PEFT baselines to fine-tune both GPT-2 Medium (355M) and GPT-2 Large (774M) models (Radford et al., 2019) on the well-established E2E (End-to-End) NLG Challenge benchmark (Novikova et al., 2017), which focuses on restaurant domain information. The statistics of the E2E NLG Challenge benchmark and the evaluation metrics applied are detailed in C.2.
Implementation Details. We adopt the experimental configurations delineated in (Hu et al., 2021; Gu et al., 2024) for the fine-tuning of query and value weights within each Transformer block, setting a rank of $r=4$. The AdamW optimizer is employed, paired with a linear learning rate schedule over 5 epochs. The reported results represent the mean outcomes from 3 runs, each initialized with a distinct random seed, selecting the performance at the last epoch of each run for comparison. For further details on the hyperparameters utilized, refer to E.2.
Main Results.
Model(Method) # Trainable Parameters BLEU NIST METEOR ROUGE-L CIDEr
[…]
F.3 Components Ablation Study on SST-2, QNLI, and STS-B
@@
F.4 Rank $k$ of Knowledge-Based SVD Truncation
@@
F.5 Rank $r$ of Knowledge-Aware Singular-Value Adaptation
@@
F.6 Parameter Initialization of $\Delta\mathbf{W}=\Delta\mathbf{U}\Delta\mathbf{\Sigma}\Delta\mathbf{V^{\top}}$
@@
F.7 Singular-Value and Orthogonal Regularization
@@
F.8 Hyperparameter Sensitivity Analysis
@@
F.9 Efficiency and Complexity Analysis
@@ G.1 Initialization Dilemmas of $\Delta\mathbf{W}$ in PiSSA and MiLoRA
$\Delta\mathbf{W}=\mathbf{A}\mathbf{B}=\mathbf{U}_{pri}\mathbf{\Sigma}_{pri}{\mathbf{V}}_{pri}^{\top},\ \Delta\mathbf{W}\leftarrow\mathbf{W}_{pri}$ (19)
@@
$\Delta\mathbf{W}$ $=\mathbf{A^{\prime}}\mathbf{B^{\prime}}=\mathbf{U^{\prime}}_{res}\mathbf{\Sigma^{\prime}}_{res}{\mathbf{V^{\prime}}}_{res}^{\top},\ \Delta\mathbf{W}\leftarrow\mathbf{W^{\prime}}_{res}$ (23)